# Low-resolution descriptions of model neural activity reveal hidden features and underlying system properties

### v1

@@ III Results and Discussion
III.1 Matrix reconstruction and bias detection for an $N=10$ Hopfield model
We begin to examine the properties of the Hopfield model through the lenses of mapping entropy considering a small network of $N=10$ at $T=0$, with an amount of stored patterns in the range $p\in\{1,..,5\}$. The relatively small total number of reduced representations (irrespectively of the resolution level $n_{cg}$) that can be designed to describe a network of this size, $n_{map}=2^{10}-1=1023$, allows us to perform an exhaustive exploration of all of them for every value of $p$ investigated; the obtained results will thus constitute a benchmark for the following analysis on a larger system discussed in Sec. III.2.
The mapping entropy and resolution associated with each CG representation of the system were calculated as detailed in Sec. II.2, providing, at fixed number of stored patterns, a point in the corresponding $S_{map}\ vs.\ H_{S}$ plane; combined together, these results fully characterise the landscape of information loss attained while observing a simple, yet nontrivial Hopfield network in terms of a subset of its constituent units.
The $S_{map}\ vs.\ H_{S}$ graphs arising from this analysis are presented in Fig. 3, where we report results separately for each value of $p$ and further label CG mappings according to their degree of coarse-graining $n_{cg}$. Overall, two features can be readily appreciated: first, and as expected, the mapping entropy on average decreases as the number of retained neurons increases, vanishing for $n_{cg}=N$; this is due to the fact that the larger the number of elements considered in the reduced description, the smaller the amount of information that is lost on the statistical properties of the high-resolution reference.
@@ III.1 Matrix reconstruction and bias detection for an $N=10$ Hopfield model
In Eq. 12, $H_{S}^{\phi}=-\sum_{\phi}p(\phi)\ln p(\phi)$ is the fine-grained resolution of the reference model, a constant quantity that is independent of the CG mapping. On the contrary, $H_{S}$ is strictly related to the choice of the reduced representation, hence contributing to variations in $S_{map}$. We note that this in principle also holds for the third factor in Eq. 12, in which the logarithm of the degeneracy $\Omega(\Psi)$ of the CG labels is averaged over the CG probability distribution; for a sufficiently large sample of high-resolution configurations, however, this term only depends on the mapping via the amount of retained sites $n_{cg}$, and is thus constant at a fixed degree of coarse-graining HGP . In such limit, Eq. 12 thus clarifies the linear dependence of $S_{map}$ on $H_{S}$ observed in Fig. 3, according to which the lower the coarse-grained resolution of the reduced representation, the lower its information loss.
At the same time, by further analysing the behaviour of the mapping entropy as a function of the resolution at a fixed degree of coarse-graining, a closer inspection of Fig. 3 reveals that, despite the overall linearity of $S_{map}$ in $H_{S}$, for several values of $n_{cg}$ the reduced representations of the system split into distinct clusters separated by gaps in mapping entropy. This feature is particularly evident in the case of $p=2$ and $p=3$ (resp. Fig. 3b and 3c), while it is almost absent for the other amounts of stored patterns. One is then naturally led to ask why such a “classification” of CG mappings appears; critically, the reason for this is to be found in the structure of the synaptic weight matrix $J_{ij}$ characterising the interactions among the network’s constituents. To better clarify this point, we focus on the simplest case with $p=2$ retaining $n_{cg}=3$ sites—for which a zoomed version of the results presented in Fig. 3b is reported in Fig. 4; furthermore, we first consider a Hopfield network with two “biased” patterns, namely
@@
The mapping entropy analysis performed on all the possible reduced representations that can be designed for this biased Hopfield network by retaining $n_{cg}=3$ neurons is presented Fig. 4c. Also in this case we observe the appearance of two clusters separated by a gap in $S_{map}$; by investigating their composition, it emerges that the CG representations that enter the group of lower $S_{map}$ are the ones that only contain spins coming both from either the first or the second blocks in Eq. 14, thus consisting of retained neurons that are maximally interacting among themselves and decoupled with the reminder. Constructing “mixed” CG mappings that gather interacting as well as noninteracting spins results in a higher information loss on the properties of the original network. This result, obtained with a model crafted ad hoc, is consistently found in the $p=2$, random patterns model of Fig. 3b, see Fig. 4b. In fact, also in this case the mappings belonging to the lower $S_{map}$ cluster are those retaining a subset of spins whose memory states are either identical or are opposite in sign between the two patterns.
@@
Let us now investigate more in detail the behaviour of mapping entropy for a varying number of retained neurons in the low-resolution description of the network; in particular, we focus on the specific set of mappings that minimises the mapping entropy at any given value of $n_{cg}$, representing the maximally informative reduced representations that can be designed for the $N=10$ Hopfield network at each degree of coarse-graining. The results of this analysis are presented in Fig. 5, and highlight that although, as previously discussed, $S_{map}$ decreases for increasing $n_{cg}$, such decrease is interrupted by some “flattenings” that occur between pairs of consecutive values $(\bar{n}_{cg},\bar{n}_{cg}+1)$. This behavior is more evident for $p=2$ and $p=3$, reduces to a softer change of slope for $p=4$ and vanish at $p=1\ ,5$. Critically, such “steps” in $S_{map}$ suggest that adding a neuron to a maximally informative mapping with $n_{cg}=\bar{n}_{cg}$ retained neurons allows little or no further information gain about the system; the retained spin set minimizing the mapping entropy at fixed $n_{cg}=\bar{n}_{cg}$ thus stands out from the others. This behaviour is comparable to the one observed by Giulini and coworkers HGP for a discrete system of non-interacting spins, where they observed an entire range of $n_{cg}$ values featuring almost the same mapping entropy minimum. In that case, this behaviour could be explained by noting that the mappings minimising $S_{map}$ most frequently included spins whose probability to be in a given state (e.g. $+1$) was appreciably different from that of the others, the latter being prone to be treated as noise. Here we obtained a similar result, despite the crucial difference of dealing with a strongly interacting system.
@@
III.2 CG’ing regimes for an $N=100$ Hopfield model
@@
The larger the size of a neural network, the richer and more complex its behaviour, which is at least in part the reason behind the emergent properties of our brain MBF . Dealing with a finite but large number of neurons $N$, however, makes the network investigation through the MEOW approach more difficult; in fact, the number of possible mappings that we have to take into account when performing decimation grows like $2^{N}$, preventing us from carrying out an exhaustive enumeration of all possible coarse-grained representations of the system. Nonetheless, resuming the discussion in the previous section, what we are interested in is not an extensive analysis of the whole CG mapping space, but rather the properties of the mappings that minimise the mapping entropy as a function of resolution.
@@
For $2\leq p\leq 8$, starting from the coarser representations, the information loss decreases with an approximately constant rate until it reaches a sort of “inflection region”. Here, the steepness of the $S_{map}\ vs.\ H_{S}$ curve decreases considerably before it grows again, and the mapping entropy continues its descent. The inflection region includes CG mappings relative to two to four values of $n_{cg}$, and shows a smearing of these mappings on a broad interval of resolutions. We thus observe a resolution gap between the two regions where the mapping entropy decreases at a constant rate.
@@
Figure 9: Resolutions of the CG representations visited during SA runs at different values of $n_{cg}$ and $p=4$ (panel a) and $p=5$ (panel b). The various resolutions explored in one SA process at given $n_{cg}$ and $p$ have the same color, and the ones corresponding to the $S_{map}$ minima found by the algorithm are indicated by black circles. The results for the different runs at fixed $p$ are plotted side by side to make it easier to compare the resolutions at which information loss is minimized for varying numbers of retained neurons.
@@
Another informative way of reading the $S_{map}-H_{S}$ graphs is to inspect the low $S_{map}$ mappings starting from those with low $n_{cg}$ (on the left of the inflection region) and studying their properties as the resolution is increased. As we pointed out for the model with $N=10$, the presence of a sudden decrease in the information loss rate suggests that increasing the detail of our CG description does not provide any significant gain in information about the reference system. The minimisation of the mapping entropy is then calling our attention to these optimal mappings, which we can investigate by looking at their values of synaptic matrix semi-dispersion, see Eq. 15.
@@
Fig. 10 displays, for each value of $p$ under examination, the semi-dispersion distributions of the CG mappings that minimize the mapping entropy. We intentionally distinguished between mappings relative to resolutions lower than, included in, or higher than the inflection region. This highlights once again a gap that separates mappings before and after this region: low-resolution mappings (in light grey) present high values of semi-dispersion, while the latter is low or even negative for high-resolution mappings (in dark grey). The CG representations that form the inflection region (shown in green in the plots) take instead intermediate semi-dispersion values, signaling a transition between these two regimes.
@@
We thus have two distinct regimes of CG representations, corresponding to different values of the resolution as well as the semi-dispersion, separated by a third, crossover interval. CG mappings tend to fall either in one regime or the other depending on their number of retained neurons, with the exception of a few specific values of $n_{cg}$ that constitute the intermediate phase where the distribution is wider and covers a large range of semi-dispersion values.
@@
The first regime includes all the mappings with lower numbers of retained neurons: in this case, the mapping entropy minimization favours CG representations with high values of semi-dispersion and high compression, which means that the most informative way of describing the reference model is to divide it into almost disconnected groups. Every highly compressed informative mapping constitutes a block of strongly interacting neurons that are effectively decoupled from the rest of the network, the dynamics of the latter being seen as an effective noise. The more neurons we retain in our description, the fewer strongly coupled spins are present and the more their dynamics will depend on the discarded neurons, until this strategy eventually becomes inadequate.
@@
The second regime shows us an alternative strategy, which includes all the mappings with higher numbers of retained neurons: when their number becomes too large for a neat decoupling of their group from the rest, the most informative description of the model features null or negative values of semi-dispersion; this indicates that the optimal representation involves weakly interacting neurons, which are often tightly coupled with the discarded ones. When this external interactions are strong, the statistical behaviour of the discarded spins is tightly bound to that of the retained ones; hence, high-resolution, informative mappings are seemingly able to describe the whole network state with the greatest detail, enabling the distinction between the different retrieved patterns.
@@ IV Conclusions
Moving to Hopfield models with $N=100$ neurons, we showed that the minima of the mapping entropy feature a regular, linearly decreasing trend as a function of resolution, with the exception of a relatively flat region where the value of $S_{map}$ remains constant for a given number $n_{cg}$ of retained neurons. This behaviour allowed us to distinguish three regions and, correspondingly, three qualitatively distinct behaviours.
@@
The first regime pertains to the most informative CG mappings with the lowest resolutions, or, equivalently, with the lowest $n_{cg}$. These mappings correspond to some of the most compressed representations that can be obtained at those values of $n_{cg}$ and their related semi-dispersion takes positive values. The conclusion is that, for low numbers of retained neurons, the mapping entropy selects those representations that describe specific clusters of neurons with strong internal interactions; the latter can thus be effectively decoupled from the rest of the network, which is treated as effective noise. This is the same coarse-graining mechanism encountered for the $N=10$ model.
@@
A rather opposite outcome is obtained for the second CG’ing regime, where we find the most informative CG mappings with higher resolutions and higher $n_{cg}$. These mappings correspond to some of the least compressed representations that can be obtained at those $n_{cg}$, and their related semi-dispersion values are low or even negative. For higher numbers of retained neurons, the mapping entropy optimisation workflow selects those representations that contain spins having weak couplings between each other, and strong couplings with the discarded ones; the optimal description of the entire network state is thus obtained retaining those that strongly correlate with the neglected ones.
@@
The third CG regime, whereby the mapping entropy remains constant, is made up of equally informative mappings that, however, vary appreciably in terms of resolution as well as semi-dispersion. These representations entail the same amount of information in spite of the fact that their “coarse-graining strategy” shifts continuously from the first to the second; such a degeneracy indicates the existence of an “information plateau”, a buffer region where varying the representations of the system does not appreciably increases or decreases its informativeness.
@@
In conclusion, this work has shown that it is possible to gather useful knowledge about a model neural network through the analysis of the information content of its low-resolution representations; in particular, we have seen that the elements of the network that are pinpointed as the most informative ones depend on the specific coarse-grained resolution level at which the system is described: this result can serve as a guide in the study of complex systems as well as in the construction of effective models of the latter, and paves the way to the development of a semi-automated protocol for the analysis of limited data gathered from systems composed by a large number of constituents.

### v2

@@ III Results and Discussion
III.1 Matrix reconstruction and bias detection for a N=10 Hopfield model
In this section we report the results of the analysis of a relatively small network, with the aim of building an intuition of the properties of the mapping entropy of optimal low-resolution representations in the context of a fully-controllable system. Specifically, the key feature of this model is the viability of exhaustively enumerating all possible decimated mappings one can employ to study the network. Hence, we can identify the absolute minima of $S_{map}$ as a function of the number of retained neurons and inspect the corresponding mappings. Furthermore, we illustrate a strategy to approximately reconstruct the interaction matrix underlying the model from an analysis of the optimal decimated representations.
We thus begin to examine the properties of the Hopfield model through the lenses of mapping entropy considering a small network of $N=10$ at $T=0$, with an amount of stored patterns in the range $p\in\{1,..,5\}$. As anticipated, the relatively small total number of reduced representations (irrespectively of the resolution level $n_{cg}$) that can be designed to describe a network of this size, $n_{map}=2^{10}-1=1023$, allows us to perform an exhaustive exploration of all of them for every value of $p$ investigated; the state configurations sampled by each setup in the course of its simulation were provided as input to the EXCOGITO software for the MEOW analysis. The obtained results will thus constitute a benchmark for the following analysis on a larger system discussed in Sec. III.2.
III.1.1 Properties of the minima of the mapping entropy and their relation with the memory patterns
The mapping entropy and resolution associated with each reduced representation of the system were calculated as detailed in Sec. II.2, providing, at fixed number of stored patterns, a point in the corresponding $S_{map}\ vs.\ H_{S}$ plane; combined together, these results fully characterise the landscape of information loss attained while observing a simple, yet nontrivial Hopfield network in terms of a subset of its constituent units.
The $S_{map}\ vs.\ H_{S}$ graphs arising from this analysis are presented in Fig. 3, where we report results separately for each value of $p$ and further label decimation mappings according to their degree of detail $n_{cg}$. Overall, two features can be readily appreciated: first, and as expected, the mapping entropy on average decreases as the number of retained neurons increases, vanishing for $n_{cg}=N$; this is due to the fact that the larger the number of elements considered in the reduced description, the smaller the amount of information that is lost on the statistical properties of the high-resolution reference.
@@ III.1.1 Properties of the minima of the mapping entropy and their relation with the memory patterns
In Eq. 12, $H_{S}^{\phi}=-\sum_{\phi}p(\phi)\ln p(\phi)$ is the fine-grained resolution of the reference model, a constant quantity that is independent of the mapping. On the contrary, $H_{S}$ is strictly related to the choice of the reduced representation, hence contributing to variations in $S_{map}$. We note that this in principle also holds for the third factor in Eq. 12, in which the logarithm of the degeneracy $\Omega(\Psi)$ of the decimated representation labels is averaged over the corresponding probability distribution; for a sufficiently large sample of high-resolution configurations, however, this term only depends on the mapping via the amount of retained sites $n_{cg}$, and is thus constant at a fixed degree of detail HGP . In such limit, Eq. 12 thus clarifies the linear dependence of $S_{map}$ on $H_{S}$ for fixed value of $n_{cg}$ observed in Fig. 3, according to which the lower the resolution of the reduced representation, the lower its information loss.
At the same time, by further analysing the behaviour of the mapping entropy as a function of the resolution at a fixed degree of detail, a closer inspection of Fig. 3 reveals that, despite the overall linearity of $S_{map}$ in $H_{S}$, for several values of $n_{cg}$ the reduced representations of the system split into distinct clusters separated by gaps in mapping entropy. This feature is particularly evident in the case of $p=2$ and $p=3$ (resp. Fig. 3b and 3c), while it is almost absent for the other amounts of stored patterns. One is then naturally led to ask why such a “classification” of decimation mappings appears; critically, the reason for this is to be found in the structure of the synaptic weight matrix $J_{ij}$ characterising the interactions among the network’s constituents. To better clarify this point, we focus on the simplest case with $p=2$ retaining $n_{cg}=3$ sites—for which a zoomed version of the results presented in Fig. 3b is reported in Fig. 4; furthermore, we first consider a Hopfield network with two “biased” patterns, namely
@@
The mapping entropy analysis performed on all the possible reduced representations that can be designed for this biased Hopfield network by retaining $n_{cg}=3$ neurons is presented Fig. 4c. Also in this case we observe the appearance of two clusters separated by a gap in $S_{map}$; by investigating their composition, it emerges that the low-resolution representations that enter the group of lower $S_{map}$ are the ones that only contain spins coming both from either the first or the second blocks in Eq. 14, thus consisting of retained neurons that are maximally interacting among themselves and decoupled with the reminder. Constructing “mixed” decimation mappings that gather interacting as well as noninteracting spins results in a higher information loss on the properties of the original network. This result, obtained with a model crafted ad hoc, is consistently found in the $p=2$, random patterns model of Fig. 3b, see Fig. 4b. In fact, also in this case the mappings belonging to the lower $S_{map}$ cluster are those retaining a subset of spins whose memory states are either identical or are opposite in sign between the two patterns.
@@
Let us now investigate more in detail the behaviour of mapping entropy for a varying number of retained neurons in the low-resolution description of the network; in particular, we focus on the specific set of mappings that minimises the mapping entropy at any given value of $n_{cg}$, representing the maximally informative reduced representations that can be designed for the $N=10$ Hopfield network at each degree of resolution. The results of this analysis are presented in Fig. 5, and highlight that although, as previously discussed, $S_{map}$ decreases for increasing $n_{cg}$, such decrease is interrupted by some “flattenings” that occur between pairs of consecutive values $(\bar{n}_{cg},\bar{n}_{cg}+1)$. This behavior is more evident for $p=2$ and $p=3$, reduces to a softer change of slope for $p=4$ and vanish at $p=1\ ,5$. Critically, such “steps” in $S_{map}$ suggest that adding a neuron to a maximally informative mapping with $n_{cg}=\bar{n}_{cg}$ retained neurons allows little or no further information gain about the system; the retained spin set minimizing the mapping entropy at fixed $n_{cg}=\bar{n}_{cg}$ thus stands out from the others. This behaviour is comparable to the one observed by Giulini and coworkers HGP for a discrete system of non-interacting spins, where they observed an entire range of $n_{cg}$ values featuring almost the same mapping entropy minimum. In that case, this behaviour could be explained by noting that the mappings minimising $S_{map}$ most frequently included spins whose probability to be in a given state (e.g. $+1$) was appreciably different from that of the others, the latter being prone to be treated as noise. Here we obtained a similar result, despite the crucial difference of dealing with a strongly interacting system.
@@
III.1.2 Approximate reconstruction of the interaction matrix from the optimal mappings
@@
III.2 Decimation regimes for a N=100 Hopfield model
@@
In this section we push forward the analysis carried out in the previous paragraphs, through the application of the MEOW approach to a Hopfield network constituted by a greater number of spins. We first report on the strategies implemented to overcome the challenges that a larger network presents to the identification of the mapping entropy minima; then, we show that this more complex system displays particularly interesting characteristics, regarding in particular the curve of mapping entropy minima as a function of the resolution of the reduced representations; finally, we inspect the features of the interactions between the groups of spins that constitute the optimal mappings, and how they change as the resolution of the simplified model is increased.
The larger the size of a neural network, the richer and more complex its behaviour, which is at least in part the reason behind the emergent properties of our brain MBF . Dealing with a finite but large number of neurons $N$, however, makes the network investigation through the MEOW approach more difficult; in fact, the number of possible mappings that we have to take into account when performing decimation grows like $2^{N}$, preventing us from carrying out an exhaustive enumeration of all possible coarse-grained representations of the system. Nonetheless, resuming the discussion in the previous section, what we are interested in is not an extensive analysis of the whole mapping space, but rather the properties of the mappings that minimise the mapping entropy as a function of resolution.
To this end, we implemented the mapping entropy optimisation workflow relying on a simulated annealing (SA) minimisation strategy ref_62_JCTC ; ref_63_JCTC , as it was originally done by Giulini et al. in Ref. JCTC . We performed simulations of an $N=100$ Hopfield network at $T=0.2$ for different values of $p\in$2, 3, …, 10}; these sets of parameters were chosen to fall within the retrieval phase of the infinite-size model, see Sec. II.1. We then looked for the maximally informative low-resolution representations of the system, i.e. the ones that minimise the mapping entropy at fixed $n_{cg}\in$1, 2, …, 20}.
III.2.1 Analysis of the least mapping entropy curve as a function of the resolution of reduced representations
@@
For $2\leq p\leq 8$, starting from the coarser representations, the information loss decreases with an approximately constant rate until it reaches a sort of “inflection region”. Here, the steepness of the $S_{map}\ vs.\ H_{S}$ curve decreases considerably before it grows again, and the mapping entropy continues its descent. The inflection region includes mappings relative to two to four values of $n_{cg}$, and shows a smearing of these mappings on a broad interval of resolutions. We thus observe a resolution gap between the two regions where the mapping entropy decreases at a constant rate.
@@
Figure 9: Resolutions of the reduced representations visited during SA runs at different values of $n_{cg}$ and $p=4$ (panel a) and $p=5$ (panel b). The various resolutions explored in one SA process at given $n_{cg}$ and $p$ have the same color, and the ones corresponding to the $S_{map}$ minima found by the algorithm are indicated by black circles. The results for the different runs at fixed $p$ are plotted side by side to make it easier to compare the resolutions at which information loss is minimized for varying numbers of retained neurons.
@@
Another informative way of reading the $S_{map}-H_{S}$ graphs is to inspect the low $S_{map}$ mappings starting from those with low $n_{cg}$ (on the left of the inflection region) and studying their properties as the resolution is increased. As we pointed out for the model with $N=10$, the presence of a sudden decrease in the information loss rate suggests that increasing the detail of our reduced representations does not provide any significant gain in information about the reference system. The minimisation of the mapping entropy is then calling our attention to these optimal mappings, which we can investigate by looking at their values of synaptic matrix semi-dispersion, see Eq. 15.
@@
Fig. 10 displays, for each value of $p$ under examination, the semi-dispersion distributions of the mappings that minimize the mapping entropy. We intentionally distinguished between mappings relative to resolutions lower than, included in, or higher than the inflection region. This highlights once again a gap that separates mappings before and after this region: low-resolution mappings (in light grey) present high values of semi-dispersion, while the latter is low or even negative for high-resolution mappings (in dark grey). The mappings that form the inflection region (shown in green in the plots) take instead intermediate semi-dispersion values, signaling a transition between these two regimes.
@@
We thus have two distinct regimes of decimated representations, corresponding to different values of the resolution as well as the semi-dispersion, separated by a third, crossover interval. Mappings tend to fall either in one regime or the other depending on their number of retained neurons, with the exception of a few specific values of $n_{cg}$ that constitute the intermediate phase where the distribution is wider and covers a large range of semi-dispersion values.
@@
The first regime includes all the mappings with lower numbers of retained neurons: in this case, the mapping entropy minimization favours reduced representations with high values of semi-dispersion and high compression, which means that the most informative way of describing the reference model is to divide it into almost disconnected groups. Every highly compressed informative mapping constitutes a block of strongly interacting neurons that are effectively decoupled from the rest of the network, the dynamics of the latter being seen as an effective noise. The more neurons we retain in our description, the fewer strongly coupled spins are present and the more their dynamics will depend on the discarded neurons, until this strategy eventually becomes inadequate.
@@
The second regime shows us an alternative strategy, which includes all the mappings with higher numbers of retained neurons: when their number becomes too large for a neat decoupling of their group from the rest, the most informative description of the model features null or negative values of semi-dispersion; this indicates that the optimal representation involves weakly interacting neurons, which are often tightly coupled with the discarded ones. When these external interactions are strong, the statistical behaviour of the discarded spins is tightly bound to that of the retained ones; hence, high-resolution, informative mappings are seemingly able to describe the whole network state with the greatest detail, enabling the distinction between the different retrieved patterns.
@@
III.2.2 Properties of the neurons most frequently retained in the optimal mappings as function of the resolution
@@ IV Conclusions
Moving to Hopfield models with $N=100$ spins, we showed that the minima of the mapping entropy feature a regular, linearly decreasing trend as a function of resolution, with the exception of a relatively flat region where the value of $S_{map}$ remains constant for a given number $n_{cg}$ of retained neurons. This behaviour allowed us to distinguish three regions and, correspondingly, three qualitatively distinct behaviours.
@@
The first regime pertains to the most informative mappings with the lowest resolutions, or, equivalently, with the lowest $n_{cg}$. These mappings correspond to some of the most compressed representations that can be obtained at those values of $n_{cg}$ and their related semi-dispersion takes positive values. The conclusion is that, for low numbers of retained neurons, the mapping entropy selects those representations that describe specific clusters of neurons with strong internal interactions; the latter can thus be effectively decoupled from the rest of the network, which is treated as effective noise. This is the same simplified representation mechanism encountered for the $N=10$ model.
@@
A rather opposite outcome is obtained for the second decimation regime, where we find the most informative mappings with higher resolutions and higher $n_{cg}$. These mappings correspond to some of the least compressed representations that can be obtained at those $n_{cg}$, and their related semi-dispersion values are low or even negative. For higher numbers of retained neurons, the mapping entropy optimisation workflow selects those representations that contain spins having weak couplings between each other, and strong couplings with the discarded ones; the optimal description of the entire network state is thus obtained retaining those that strongly correlate with the neglected ones.
@@
The third regime, whereby the mapping entropy remains constant, is made up of equally informative mappings that, however, vary appreciably in terms of resolution as well as semi-dispersion. These representations entail the same amount of information in spite of the fact that their “decimation strategy” shifts continuously from the first to the second; such a degeneracy indicates the existence of an “information plateau”, a buffer region where varying the representations of the system does not appreciably increases or decreases its informativeness.
The results presented in this work have highlighted nontrivial relations between the level of detail of the low-resolution representation at which the system is inspected and the underlying generative process, specifically the interaction matrix of the model. The analysis has focused on particular realisation of the Hopfield network, with the aim of establishing a link the most direct possible between a case-specific system and its optimal low-resolution mappings. The picture that emerges thus provides promising evidence that such link can be leveraged to extract nontrivial information about the properties of systems when only part of them is accessible to inspection. Further work is required to assess the generality of the observed features; in particular, a broad-range campaign of simulations and analysis is in order to characterise the typical expected behaviour in various setups, e.g., averaging over the quenched disorder for different degrees of structure in the statistical properties of the stored patterns.
@@
In conclusion, this work has shown that it is possible to gather useful knowledge about a model neural network through the analysis of the information content of its low-resolution representations; in particular, we have seen that the elements of the network that are pinpointed as the most informative ones depend on the specific decimated representation resolution level at which the system is described: this result can serve as a guide in the study of complex systems as well as in the construction of effective models of the latter, and paves the way to the development of a semi-automated protocol for the analysis of limited data gathered from systems composed by a large number of constituents.